OpenStudy (anonymous):

how do I find the LCD of 1/6, 1/10, 1/18 and 1/21?

OpenStudy (mertsj):

Write the prime factors of all denominators in exponent form and then select and multiply the largest power of every DIFFERENT factor.

OpenStudy (amistre64):

1/6, 1/10, 1/18 and 1/21 = lcd 1, 6/10, 6/18 and 6/21 = 6lcd 1, 3/5, 1/3 and 2/7 = 6lcd 3, 9/5, 1 and 6/7 = 18lcd 15, 9, 5 and 30/7 = 90lcd 105, 63, 35 and 30 = 630 lcd maybe

OpenStudy (mertsj):